OpenStudy (anonymous):

Find the exact value given the following cos theta =2/9 and theta in quadrant IV find sin theta Help plzzzz

OpenStudy (anonymous):

\[\sqrt{c ^2-a^2}=b\]

OpenStudy (anonymous):

\[\sqrt{2^2-9^2}=\]

OpenStudy (anonymous):

\[\sqrt{4-81}= 8.02\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Depends on if you're teacher wants it in decimal form or not. Since it says exact value its probably best to leave it as \[\sqrt{77}\]

OpenStudy (anonymous):

Oh and, since it's in quadrant 4, and we are looking for the y value, it's negative

OpenStudy (anonymous):



OpenStudy (anonymous):

Sin = opposite/hypotenuse So sin = \[\frac{ -\sqrt{77} }{ 9 }\]